OpenStudy (andriod09):

How do you solve: \[\frac{7-5x}{4}=-2\]

OpenStudy (gorica):

multiply both sides by 4, you get 7-5x=-8 7+8=5x 15=5x x=15/3 x=3

OpenStudy (gorica):

* x=15/5

OpenStudy (andriod09):

I'm sorry, but that makes no sense. I don't know how to get the 7 down from the \[\frac{7-5x}{4}\] if i could do that, then i know how to solve the problem.

OpenStudy (gorica):

\[\frac{ 7-5x }{ 4 } 4 = 7-5x\]

OpenStudy (andriod09):

Wheres the reciprocal then?

OpenStudy (andriod09):

because you can't do \[\frac{4}{7-5x}\] with out messing up the equation, and you can't just multiply them by 4.

OpenStudy (gorica):

you can multiply both sides of equation by the same number and you won't change the result. So, multiply fraction by 4 and multiply -2 by 4 and you will get what I wrote above.

OpenStudy (andriod09):

But that's not right. If you had a problem like: \[\frac{9}{4}k=-36\] you would do this: \[(\frac{4}{9})(\frac{9}{4})k=-36(\frac{4}{9})\] then you would divide -36 by 9, to get four, then four times four. How would you do that to this problem?

OpenStudy (gorica):

if you multiply \[\frac{ 9 }{ 4 }k=-36\] by 4, you'll get 9k=-36*4 then you divide this equation by 9 and you get k=-\[k=-\frac{ 36 }{ 9 } 4\] k=-4*4 k=-16 You will get the same result if you multiply equation from the start by 4/9. Right? Ok, if you want reciprocal for your first question, you can write 4 as \[\frac{ 4 }{ 1 }\]. Ok?

OpenStudy (andriod09):

But that's not the reciprocal! The reciprocal of the problem would cancel them out, and I can't do it like that!

OpenStudy (gorica):

ok, tell me, what is reciprocal for \[\frac{ 1 }{ 4 }\]?

OpenStudy (andriod09):

The reciprocal for \[\frac{1}{4}\] is \[\frac{4}{1}\] but the reciprocal for \[\frac{7-5x}{4}\] isn't 4, its \[\frac{4}{7-5x} which would cancel out the x, and just leave you with a normal number.

OpenStudy (gorica):

you can't cancel out x because you have to find a value of x

OpenStudy (gorica):

that's the aim, to find value of x.

OpenStudy (andriod09):

That's what i'm saying!

OpenStudy (gorica):

Do you know that \[\frac{ a }{ b }=\frac{ c }{ d }\] is the same as a*d=c*b ?

OpenStudy (andriod09):

Yes i do, But you're seriously confusing me.

OpenStudy (gorica):

Ok, good! Then let me help you.

OpenStudy (andriod09):

Then make it easily understandable!

OpenStudy (gorica):

\[\frac{ 7-5x }{ 4 }=-2\] is the same as \[\frac{ 7-5x }{ 4 }=\frac{ -2 }{ 1 }\] because \[-2=\frac{ -2 }{ 1 }\] now you have, from \[\frac{ a }{ b }=\frac{ c }{ d } <=> a*d=c*b\], (7-5x)*1=4*(-2) 7-5x=-8 7+8=5x 15=5x x=3

OpenStudy (gorica):

ok?

OpenStudy (phi):

you misunderstand the "multiply by the reciprocal" rule in your problem \[ \frac{1}{4} \cdot (7-5x) = -2\] one way to solve is to multiply both sides by 4 \[ \frac{4}{4} \cdot (7-5x) = -2\cdot 4\]

OpenStudy (andriod09):

That's not my problem though, my problem is this: \[\frac{7-5x}{4}\] that's what it is.

OpenStudy (phi):

as you know, \[ \frac{7-5x}{4} = \frac{1}{4} \cdot (7-5x) \]

OpenStudy (gorica):

I'm giving up :D

OpenStudy (phi):

another way to write it is \[ \frac{7-5x}{4} = \frac{7}{4} - \frac{5x}{4} \]

OpenStudy (phi):

dividing by 4 is the same as multiplying by 1/4

OpenStudy (andriod09):

I don't get it.

OpenStudy (phi):

which part ?

OpenStudy (andriod09):

The reciprocal part. How do you get 1/4??

OpenStudy (phi):

maybe it makes more sense if we use numbers as examples: Example: \[ \frac{8}{4}=2 \\ \text{ and } \\ \frac{1}{4} \cdot 8= 2\] dividing 8 by 4 is the same as multiplying 8 by 1/4

OpenStudy (andriod09):

So how do you get 1/4 from 7-5x/4? that's what i need to know...

OpenStudy (phi):

if you know how to multiply fractions (top times top over bottom times bottom) we can say \[ \frac{1}{4} \cdot 8 = \frac{1}{4} \cdot\frac{8}{1}=\frac{1 \cdot 8}{4 \cdot 1}= \frac{8}{4} = 2 \]

OpenStudy (phi):

that same rule works for complicated expressions like 7-5x (instead of 8)

OpenStudy (andriod09):

OKay, but still, how do you separate the 7/4 from the -5x/4?

OpenStudy (phi):

Here is an example of how to add two fractions with the same denominator, \[ \frac{ 1}{4} + \frac{3}{4} \] the answer is : keep the same denominator and add the numerators \[ \frac{ 1}{4} + \frac{3}{4}= \frac{1+3}{4}\] notice you can switch between those two forms example \[ \frac{7-5x}{4} = \frac{7}{4} + \frac{-5x}{4} \]

OpenStudy (phi):

another way to solve your problem (though this is a bit messy) \[ \frac{7}{4} + \frac{-5x}{4}= -2 \\ \frac{-5x}{4}= -2-\frac{7}{4} \] simplify, then multiply both sides by -4/5

OpenStudy (andriod09):

okay, i'm beyond confused. >.< i just need to answer these last few questions...

OpenStudy (phi):

can you solve for x in \[ \frac{4}{4} \cdot (7-5x) = -2\cdot 4 \]?

OpenStudy (phi):

4/4 is 1 -2*4 is -8 so you get \[ 7 -5x= -8\] can you finish ?

OpenStudy (andriod09):

\[7-5x=-8\]\[-7=-7\]\[-5x=15\]\[\frac{-5x}{5}=\frac{-15}{-5}\]\[x=3\] right?

OpenStudy (phi):

yes, though you have a typo on the 2nd to last line: you mean -5x/ -5 but that is the answer. Now to do the original problem you have to convince your self that multiplying both sides by 4 is a good thing to do.

OpenStudy (andriod09):

Okay, but still, how would you get the four from \[\frac{7-5x}{4} That's what i don't understand.

OpenStudy (phi):

Can you answer this \[ 4 \cdot \frac{x}{4} = \]?

OpenStudy (andriod09):

no. not correctly at least.

OpenStudy (phi):

can you do \[ 4 \cdot \frac{2}{4} \]?

OpenStudy (andriod09):

uhh. would it be: \[\frac{8}{16}\] or \[\frac{1}{2}\]?

OpenStudy (phi):

a calculator would tell you. But if you don't understand fractions you will get stuck doing algebra. when you get time, see http://www.khanacademy.org/math/arithmetic/fractions/multiplying_and_dividing_frac/v/multiplying-fractions (and any of the videos about fractions that you don't know) but the rule for multiplying fractions is multiply top times top and bottom times bottom if one of the numbers is a whole number, set its "bottom" to 1 \[ 4 \cdot \frac{2}{4}= \frac{4}{1}\cdot \frac{2}{4} = \frac{4\cdot 2}{1\cdot 4}\]

OpenStudy (phi):

one of the rules you know about multiplying is you can swap the order. so \[ \frac{4\cdot 2}{1\cdot 4} = \frac{2\cdot 4}{1\cdot 4}\] we can think of that as multiplying \[ \frac{2}{1}\cdot \frac{4}{4}\] of course 4/4 is 1 \[ \frac{2}{1}\cdot 1 \] and that simplifies to 2 if you practice, you will learn that if you have the same thing in the top and bottom (like the 4 in this example), you can cancel it. and get 2 immediately

OpenStudy (phi):

so if you see \[ 4 \cdot \frac{2}{4}= \cancel{4} \cdot \frac{2}{\cancel{4}} =2 \]

OpenStudy (phi):

\[ \frac{7-5x}{4} = -2 \\ 4 \cdot \frac{7-5x}{4} = -2 \cdot 4 \] now use the rule "cancel" the 4, because you have a 4 in the top and bottom